Find a function of minimal gegree with interger coefficients and solutions { -4, 3/2, 2+i }. Leave answer in factored form.
Answer by ankor@dixie-net.com(22740)

Find a function of minimal gegree with interger coefficients and solutions { -4, 3/2, 2+i }. Leave answer in factored form.
:
x = -4
(x+4)
:
x = 3/2
(2x-3)
:
x = 2 + i
x - 2 = i
Square both sides
(x-2)^2 = i^2
x^2 - 4x + 4 = -1
x^2 -4x + 4 + 1 = 0
(x^2 - 4x + 5)
;
:
Three factors: (x + 4)(2x - 3) (x^2 - 4x + 5)